OpenStudy (goldeverything):

Find the solution set of x. x plus thirty-four point zero one is less than forty-four point eight x is less than ten point seven nine x is greater than ten point seven nine x is less than seventy eight point eight one x is greater than seventy eight point eight one MEDAL WILL BE AWARDED

OpenStudy (anonymous):

x+34.01<41.8 x<34.01-41.8 = ...

OpenStudy (anonymous):

Correction... x <41.8-34.01 = ...

OpenStudy (goldeverything):

so what is the awnser A,B,C,D

OpenStudy (anonymous):

I mistyped...again x < 44.8-34.01 = 10.79

OpenStudy (goldeverything):

thank you